Bob has 60 coins consisting of quarters and dimes. The coins combined value is $9.45. Find out how many are quarters and how many are dimes
Answer by stanbon(75887)

Bob has 60 coins consisting of quarters and dimes. The coins combined value is $9.45. Find out how many are quarters and how many are dimes
-------
Quantity Eq:: d + q = 60 coins
Value Eq:: 10d+25q = 945 cents
-------------------------------------
Modify:
2d + 2q = 120
2d + 5q = 189
----------------
Subtract:
3q = 69
q = 23 (# of quarters)
---
Solve for "d":
d+q = 60
d + 23 = 60
d = 37 (# of dimes)
========================
Cheers,
stan H.